17. The two sets of grandparents for a newborn baby wish to invest enough money immediately to pay $10,000 per year for four years toward college costs starting at age 18. Grandparents A agree to fund the first two payments, while Grandparents B agree to fund the last two payments. If the effective rate of interest is 6% per annum. find the difference between the contributions of Grandparents A and B.
Answer #1

a. Present Value of First set of Grandparents Payments

Present Value = Amount * (1 / 1.06^18) + Amount * (1 / 1.06^19)

Present Value = 10000 * (1 / 1.06^18) + 10000 * (1 / 1.06^19)

Present Value = 10000 * 0.350343 + 10000 * 0.330513

Present Value = $6808.57

b. Present Value of Second set of Grandparents Payments

Present Value = Amount * (1 / 1.06^18) + Amount * (1 / 1.06^19)

Present Value = 10000 * (1 / 1.06^20) + 10000 * (1 / 1.06^21)

Present Value = 10000 * 0.311805 + 10000 * 0.294155

Present Value = $6059.60

c. Difference between contribution = $6808.57 - $6059.60 = $748.97